OpenStudy (anonymous):

A rescue helicopter flew from its home base for 35 km on a course of 330 degrees to pick up an accident victim. It then flew 25 km on a course of 090 degrees to the hospital. What distance and on what course will the helicopter fly to directly return to its home base?

OpenStudy (anonymous):

330 degrees is the same as 30 degrees to the southeast. 35cos(30 degrees) km east = 30.3108891324553 km east 35sin(30 degrees) km south = 17.5 km south Next he flew 25 km due north So overall, he flew 30.3108891324553 km east, and (25 - 17.5) = 7.5 km north So he must fly southwest to get back. Fly south 7.5 km, and Fly west 30.3108891324553 km Magnitude of distance is sqrt(7.5^2 + 30.3108891324553 ^2) = 31.225 km (looks good so far) Direction = 180 + arctan(7.5/30.3108891324553 ) = 194 degrees

OpenStudy (anonymous):

I'm looking at a unit circle and are you saying that 330 degrees southeast = 30 degrees northeast?

OpenStudy (anonymous):

how would it look like? like this: 

OpenStudy (anonymous):

can someone please draw it out? I don't understand..

OpenStudy (stormfire1):

Since this is a question involving navigation, you must use compass degrees...which is different than what a cartesian plane would show. Here's a picture which should help...you just need to solve the triangle to find the length of the of last line (and angle) for the plane to get back to home base.